OpenStudy (anonymous):

user polar coordinates to evaluate the integral of region of (x+y)dA, region r is bounded by the graph of x^2+y^2 = 2y

OpenStudy (experimentx):

x+y is a plane

OpenStudy (anonymous):

i just don't get how to get the limits =\

OpenStudy (experimentx):

radius ... 0.5

OpenStudy (anonymous):

na the limit is 2sintheta =\

OpenStudy (experimentx):

\[ \int_{0}^{1} \int_{0 }^{+ \pi} f(r, \theta) rd\theta dr\] By looking at graph!!

OpenStudy (anonymous):

but that's not right =\ \[\int\limits_{0}^{pi/2}\int\limits_{0}^{2\sin \theta} (r \cos \theta + rsin \theta) r dr d \theta\]

OpenStudy (experimentx):

Oo ... sorry!! I think I made mistake!! r is dependent on theta

OpenStudy (anonymous):

is the answer, but i have no idea how she go that

OpenStudy (experimentx):

looks like today's not my day!! http://www.wolframalpha.com/input/?i=r+%3D+2+sin+theta \[ \int_{0}^{\pi}\int_{0}^{2\sin\theta} f(r,\theta)rdrd\theta\]

OpenStudy (experimentx):

If you look at the graph ... theta goes from 0 to pi and ... r varies as 2sin\theta, so limits for r should be 0 to 2sin\theta and also integrate r first!! now sure about pi/2

OpenStudy (anonymous):

what if i don't have the graph?

OpenStudy (anonymous):

i know i'm supposed to use what x and y equal to in polar coordinates, so like x=rcostheta y=rsintheta x^2+y^2=r^2, but i don't understand how to do this in this instance..

OpenStudy (anonymous):

wait nvm i think i figured it out haha

OpenStudy (experimentx):

?? graphing your boundary is important ... it's the first thing you must do before you integrate!! dxdy = r dr d\theta

OpenStudy (experimentx):

f(x,y) = x + y => r(cos \theta + sin\theta)

OpenStudy (anonymous):

what if i don't have a graphing calculator?

OpenStudy (anonymous):

oh you were right i was looking at the wrong problem, it's o to pi ^.^

OpenStudy (anonymous):

but yeah i don't know how to graph it without a calculator =\

OpenStudy (experimentx):

you are supposed to do it by hand ... graphing curves are taught before or at single variable calculus ...!!

OpenStudy (anonymous):

i def don't remember that lol i took it four years ago :(

OpenStudy (anonymous):

lol my final is tomorrow night, i can't spend an hour listening to a lecture =\

OpenStudy (anonymous):

i'm just going to assume that since its x and y squared, it's a circle, and the 2y means that i moves up on the y axis. anyway, how do i integrate this? :( \[\int\limits_{0}^{pi}\sin^3 \theta \cos \theta + \sin^4 \theta d \theta\]

OpenStudy (experimentx):

integrate it individually ... use substitution ... \[ \int \sin^3x \cos x dx = ... \] change sin^4x = (sin^2x)^2 = ((1- cos 2x)/2)^2 ... square it ... again change cos^2 2x ... using double formula ... LOL ... it's going to long ...

OpenStudy (experimentx):

the later part was for sin^4x

OpenStudy (anonymous):

i don't even know what you just did lol

OpenStudy (experimentx):

sorry ... i didn't get notification

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=integrate+sin%5E3x+cos+x click show steps ... to view steps

OpenStudy (experimentx):

for the other http://www.wolframalpha.com/input/?i=integrate+sin%5E4x